# FedSN: A Novel Federated Learning Framework over LEO Satellite Networks

### v4

@@ Abstract
Recently, a large number of Low Earth Orbit (LEO) satellites have been launched and deployed successfully in space. Due to multimodal sensors equipped by the LEO satellites, they serve not only for communication but also for various machine learning applications. However, the ground station (GS) may be incapable of downloading such a large volume of raw sensing data for centralized model training due to the limited contact time with LEO satellites (e.g. 5 minutes). Therefore, federated learning (FL) has emerged as the promising solution to address this problem via on-device training. Unfortunately, enabling FL on LEO satellites still face three critical challenges: i) heterogeneous computing and memory capabilities, ii) limited uplink rate, and iii) model staleness. To this end, we propose FedSN as a general FL framework to tackle the above challenges. Specifically, we first present a novel sub-structure scheme to enable heterogeneous local model training considering different computing, memory, and communication constraints on LEO satellites. Additionally, we propose a pseudo-synchronous model aggregation strategy to dynamically schedule model aggregation for compensating model staleness. Extensive experiments with real-world satellite data demonstrate that FedSN framework achieves higher accuracy, lower computing, and communication overhead than the state-of-the-art benchmarks.
@@
The significant progress made in satellite technology has resulted in a sharp reduction in the expenses associated with launching satellites. This has encouraged an enormous expansion in the number of satellites orbiting the Earth at low altitudes. Several corporations have pledged to launch thousands of Low Earth Orbit (LEO) satellites over the next few years [1]. For example, SpaceX [2] plans to launch about 42,000 LEO satellites, and different from geostationary (GEO) satellites, LEO satellite networks always construct meta-constellations for communications. More importantly, LEO satellites collect large amounts of sensor data [3, 4], such as spectrum and earth imageries, which can be leveraged to resolve many global challenges such as food safety [5], disease spread [6], and climate change [7] with the power of deep learning.
It is relatively easy to train deep learning models using plentiful data in terrestrial networks in a timely and synchronous fashion [8, 9, 10, 11, 12]. However, such conditions do not hold in the space networks, since conventional centralized learning frameworks over satellite networks are significantly hindered by the limited contact time between LEO satellites and a ground station (GS). We face a fundamental problem that datasets are often too large to download to a GS [13]. Moreover, the satellite nodes may not be willing to share the data with GSs due to commercial interests and data ownership concerns if they belong to different parties. Fortunately, federated learning (FL) [14] is a new paradigm for distributed training technology on massive participating edge devices. FL only needs a central server to coordinate all participating edge devices, and aggregate their local model parameters [15, 16, 17, 18, 19]. In this way, it does not only protect data privacy, but also reduce communication overhead between a central server and edge devices [20].
@@
We propose a unified constrained budget to map the computing and memory resources, and uplink rate into a utility value. Based on this, we design an FL sub-structure scheme to handle heterogeneous computing and memory resources and limited uplink rate for LEO satellite networks to achieve superior training performance.
@@ 2.2 Bottleneck of Uplink Communication
We leverage a real-world commercial LEO satellite communication system, Starlink to measure and study its downlink and uplink rates. Fig. LABEL:sfig:exStarlink and Fig. LABEL:sfig:gsStarlink show the experimental setup of Starlink and its GS, respectively. We use a well-known network measurement tool, Iperf [29] to test and collect downlink and uplink rates, and their cumulative distribution functions (CDF) are illustrated in Fig. LABEL:sfig:data_rate that apparently, the average downlink rate is close to 100 Mbps, and larger than the average uplink rate with 12 Mbps. In LEO satellite FL, the downlink rate is enough for model aggregation, but the limited uplink rate becomes a major bottleneck for model distribution. For example, the size of a VGG-16 is 528 MB, and hence, GS spends about 5.87 minutes to distribute a single model. The time cost of distribution is too long to deliver models to all satellites in each satellite-GS contact time, and thus fewer satellites participate in training their models.
We also evaluate local models updating failure at satellite networks. According to our Starlink measurement traces, we use Linux traffic control, tc [30] to emulate satellite networks condition. We set contact time ranging from 5 to 15 minutes [31], and packet loss rate at 100% when out of contact time immediately via tc. In this experiment, we deploy three satellites333In practical deployment, the number of satellites involved during the contact time far exceeds three [2], leading to a substantial increase in the volume of data exchanged, thus posing greater challenges for uplink model transmission. Therefore, despite larger GS being equipped with a large antenna with higher transmit power, the uplink model transmission problem still cannot be ignored., and the result of that experiment is shown in Fig. LABEL:sfig:update_fail. The ideal case assuming no uplink rate limitation has the best performance, but the other two cases lose at least one satellite to participate in training leading to worse performance. The reason is that in the FL model aggregation phase, more satellites can integrate more data into model training [32]. In short, uplink communication of LEO satellites heavily affects FL training performance in practice.
@@ 3.3 Satellite Characterization
where , represents the set of satellites in the -th orbit (i.e., the -th intra-group set), is the proportion of global model training completed by -th satellite in the -th inter-group in an orbital period, ranging from 0 to 1., denotes the computing resources allocated to FL model training for the -th satellite in the -th inter-group set (namely, the number of floating point operations (FLOPs) per second), and are local update epoch (training round) and the computing workload (in FLOPs) of global model training for a mini-batch data samples, and represent the available memory resources for the -th satellite in the -th inter-group set and the data size (in bits) of the global model, respectively.
@@
where is the average uplink rate from a GS to the -th satellite in the -th inter-group set, and and represent a period of contact time and data size (in bits) of a global model, separately. For our design, jointly considering satellites’ heterogeneous computing and memory resources as well as uplink rate, the available budget for the -th satellite in the -th inter-group set is expressed as
@@ 3.5.1 Sub-structure Customization Method (SCM)
Given the above two observations, we find that the sub-structure customization needs to be well-designed, otherwise, it may lead to deterioration rather than improvement in training performance. In view of this, we propose a flexible sub-structure customization method. We first determine basic sub-structure models based on the minimum budget constraint within satellite groups, ensuring that every satellite is capable of training these models. Then, we split a global model into multiple basic sub-structure models via channel-wise partitioning and customize different combinations of those models according to satellites’ available budgets . This eliminates occurrences of under training for basic sub-structure models, thereby effectively improving the convergence speed and test accuracy of model training. We use an example to illustrate how FedSN determines the number of basic sub-structure models in Fig. 6. The global model is split into four 0.25 basic sub-structure models, and there are three satellites in the -th inter-group set with . In this case, the GS determines that the number of basic sub-structure models are , , and , for those satellites, respectively.

### v5

# FedSN: A Federated Learning Framework over Heterogeneous LEO Satellite Networks

@@ Abstract
Recently, a large number of Low Earth Orbit (LEO) satellites have been launched and deployed successfully in space. Due to multimodal sensors equipped by the LEO satellites, they serve not only for communication but also for various machine learning applications. However, the ground station (GS) may be incapable of downloading such a large volume of raw sensing data for centralized model training due to the limited contact time with LEO satellites (e.g. 5 minutes). Therefore, federated learning (FL) has emerged as the promising solution to address this problem via on-device training. Unfortunately, enabling FL on LEO satellites still face three critical challenges: i) heterogeneous computing and memory capabilities, ii) limited downlink/uplink rate, and iii) model staleness. To this end, we propose FedSN as a general FL framework to tackle the above challenges. Specifically, we first present a novel sub-structure scheme to enable heterogeneous local model training considering different computing, memory, and communication constraints on LEO satellites. Additionally, we propose a pseudo-synchronous model aggregation strategy to dynamically schedule model aggregation for compensating model staleness. Extensive experiments with real-world satellite data demonstrate that FedSN framework achieves higher accuracy, lower computing, and communication overhead than the state-of-the-art benchmarks.
@@
The significant progress made in satellite technology has resulted in a sharp reduction in the expenses associated with launching satellites. This has encouraged an enormous expansion in the number of satellites orbiting the Earth at low altitudes. Several corporations have pledged to launch thousands of Low Earth Orbit (LEO) satellites over the next few years [1]. For example, SpaceX [2] plans to launch about 42,000 LEO satellites, and different from geostationary (GEO) satellites, LEO satellite networks always construct meta-constellations for communications. More importantly, LEO satellites collect large amounts of sensor data [3, 4, 5], such as spectrum and earth imageries, which can be leveraged to resolve many global challenges such as food safety [6], disease spread [7], and climate change [8] with the power of deep learning.
It is relatively easy to train deep learning models using plentiful data in terrestrial networks in a timely and synchronous fashion [9, 10, 11, 12, 13]. However, such conditions do not hold in the space networks, since conventional centralized learning frameworks over satellite networks are significantly hindered by the limited contact time between LEO satellites and a ground station (GS). We face a fundamental problem that datasets are often too large to download to a GS [14]. Moreover, the satellite nodes may not be willing to share the data with GSs due to commercial interests and data ownership concerns if they belong to different parties [15]. Fortunately, federated learning (FL) [16] is a new paradigm for distributed training technology on massive participating edge devices. FL only needs a central server to coordinate all participating edge devices, and aggregate their local model parameters [17, 18, 19, 20]. In this way, it does not only protect data privacy, but also reduce communication overhead between a central server and edge devices.
@@
We propose a unified constrained budget to map the computing and memory resources, and uplink rate into a utility value. Based on this, we design an FL sub-structure scheme to handle heterogeneous computing and memory resources and limited downlink/uplink rate for LEO satellite networks to achieve superior training performance.
@@ 2.2 Bottleneck of Uplink Communication
We leverage a real-world commercial LEO satellite communication system, Starlink to measure and study its downlink and uplink rates. Fig. LABEL:sfig:exStarlink and Fig. LABEL:sfig:gsStarlink show the experimental setup of Starlink and its GS, respectively. We use a well-known network measurement tool, Iperf [29] to test and collect downlink and uplink rates, and their cumulative distribution functions (CDF) are illustrated in Fig. LABEL:sfig:data_rate that apparently, the average downlink rate is close to 100 Mbps, and larger than the average uplink rate with 12 Mbps. In LEO satellite FL, the downlink rate is enough for model aggregation, but the limited uplink rate becomes a major bottleneck for model distribution. For example, the size of a VGG-16 is 528 MB, and hence, GS spends about 5.87 minutes to distribute a single model. The time cost of distribution is too long to deliver models to all satellites in each satellite-GS contact time, and thus fewer satellites participate in training their models. Although specialized GS (with multi-million US dollars cost [30, 31]) offers Gbps- and hundreds of Mbps-level rates respectively for downlink and uplink, multiple tenants need to share these rates, resulting in very limited per-tenant rates and thus severely restricting their smashed data exchanges. Moreover, recent work [32] also utilizes distributed GSs of low-cost commodity hardware to receive satellite downlink signals. Therefore, results produced by our commercial GS are representative.
We also evaluate local models updating failure at satellite networks. According to our Starlink measurement traces, we use Linux traffic control, tc [33] to emulate satellite networks condition. We set contact time ranging from 5 to 15 minutes [34], and packet loss rate at 100% when out of contact time immediately via tc. In this experiment, we deploy three satellites333In practical deployment, the number of satellites involved during the contact time far exceeds three [2], resulting in lower transmission rate, thus posing greater challenges for uplink model transmission. Therefore, despite larger GS being equipped with a large antenna with higher transmit power, the uplink model transmission problem is still challenging., and the result of that experiment is shown in Fig. LABEL:sfig:update_fail. The ideal case assuming no uplink rate limitation has the best performance, but the other two cases lose at least one satellite to participate in training leading to worse performance. The reason is that in the FL model aggregation phase, more satellites can integrate more data into model training [35]. In short, uplink communication of LEO satellites heavily affects FL training performance in practice.
@@ 3.3 Satellite Characterization
where , represents the set of satellites in the -th orbit (i.e., the -th intra-group set), is the proportion of global model training completed by -th satellite in the -th inter-group in an orbital period, ranging from 0 to 1, denotes the computing resources allocated to FL model training for the -th satellite in the -th inter-group set (namely, the number of floating point operations (FLOPs) per second), and are local update epoch (training round) and the computing workload (in FLOPs) of global model training for a mini-batch data samples, and represent the available memory resources for the -th satellite in the -th inter-group set and the data size (in bits) of the global model, respectively.
@@
where , and are the average uplink and downlink rates between GS and the -th satellite in the -th inter-group set, and and represent a period of contact time and data size (in bits) of a global model, separately. For our design, jointly considering satellites’ heterogeneous computing and memory resources as well as uplink rate, the available budget for the -th satellite in the -th inter-group set is expressed as
@@ 3.5.1 Sub-structure Customization Method (SCM)
Given the above two observations, we find that the sub-structure customization needs to be well-designed, otherwise, it may lead to deterioration rather than improvement in training performance. In view of this, we propose a flexible sub-structure customization method. We first determine basic sub-structure models based on the minimum budget constraint within satellite groups, ensuring that every satellite is capable of training these models. Then, we split a global model into multiple basic sub-structure models via channel-wise partitioning and customize different combinations of those models according to satellites’ available budgets . This eliminates occurrences of under training for basic sub-structure models, thereby effectively improving the convergence speed and test accuracy of model training. We use an example to illustrate how FedSN determines the number of basic sub-structure models in Fig. 6.
The global model is split into four 0.25 basic sub-structure models, and there are three satellites in the -th inter-group set with . In this case, the GS determines that the number of basic sub-structure models are , , and , for those satellites, respectively.